OpenStudy (anonymous):

Evaluate the function H(x) = 16 – x, when x = –4. Please help

OpenStudy (anonymous):

this is similar to my last question, but a bit different, so im confused....

OpenStudy (anonymous):

h(x)=16-x h(-4)=16-(-4) h(-4)=20